# references.bib

### v5

@@ 3.3 The role of long-range correlations between metaorders
It is known that the signs of metaorders initiated by different traders are also correlated, see [donier2015million, bucci2020co]. This may either be due to herding, or more plausibly to different traders following the same signal. As mentioned above, we assume that the sign cross-correlation $\mathbb{E}[\varepsilon(t)\varepsilon(t+\tau)]$ decays as $\Gamma(\tau_{0}/\tau)^{\gamma_{\times}}$, whereas the sizes $q$ and $q^{\prime}$ are remain independent for simplicity. [^4] One can extend the following calculations to the case where conditional size distribution of a metaorder starting at $t+\tau$, knowing that one metaorder started at $t$ is $\Psi_{q}(s^{\prime}|s,\tau)=\frac{\tau^{b}}{1+\tau^{b}}\Psi_{q}(s^{\prime})+\frac{1}{1+\tau^{b}}\frac{1}{s}F(s^{\prime}/s),$ (35) where $F(.)$ is a certain function and $b$ a new exponent. The following results are unaffected provided $b>1$. When $\Gamma=0$, there is no correlation between successive metaorders.
@@
In other words, the LMF model [lillo2005theory] that ascribes the main contribution to sign autocorrelation to long metaorders is only valid if such metaorders are not too strongly correlated between themselves, i.e. when
@@
In view of the empirical data supporting the LMF model, we stick to this assumption henceforth. In fact, one can measure $\gamma_{\times}$ directly (G. Maitrier, unpublished, see also [donier2015million]) suggesting $\gamma_{\times}\approx\gamma$.
@@ 3.4.1 Child volume distribution
As discussed in Section 3.2, we consider a log-normal distribution for the child order sizes, as defined by Eq. (28). It turns out to be a reasonable approximation of reality for large volumes, see Fig. 1, with values of $\sigma_{\ell}$ reported in the legend, around $1$ for stocks and SPMINI and $1.2$ for EUROSTOXX. A better fit of the tail of the distribution is, arguably, power-law $\propto q^{-1-\varpi}$, with $\varpi$ found to be around 2.12.12.12.12.42.42.42.4. Such a choice makes the mathematical analysis more cumbersome, and we prefer the log-normal specification for our semi-quantitative discussion of the role of volume fluctuations. Nevertheless, a power-law tail can play a role similar to the coefficient $\lambda$ in $\mu_{q}$ when it comes to the scaling of $\Sigma^{(2n)}_{a}$, crossing over to a linear behaviour when $2na\gtrsim\varpi$, see footnote 3.
@@ 3.4.2 Distribution of sign imbalances
We show in Fig. 2 the distribution of rescaled sign imbalances $I^{0}T^{-\chi}$, for different $T$ in trade time and $\chi=0.72$. The theoretical analysis performed in the previous sections predicts that such distributions should collapse when $\chi$ is chosen to be $1/\mu$, where $\mu=1+\gamma$ is related to the autocorrelation of the sign of the trades, which gives $\mu\approx 1.4$. Although not perfect, the agreement is quite reasonable, in view of the fact that $\mu$ actually depends on the size of the child order $q$, see Fig. 3. The master curve is clearly non-Gaussian, with tails that become fatter as $T$ increases, as expected from our theoretical prediction.
Figure 2: Distribution of sign imbalances $I^{0}$ as a function of the window size $T$ (measured in number of trades) on EUROSTOXX, between 2016 and 2019. After a proper rescaling by $T^{-\chi}$, with $\chi=0.72$, distributions nicely collapse onto a single master curve, consistent with findings in [patzelt2018universal]. The value of $\chi$ is not far from the theoretical prediction $\chi=1/\mu$ with $\mu=3/2$. The dotted line shows a Gaussian distribution with the same variance as red curve. As expected, the volume imbalance exhibits fat tails.
3.4.3 Sign autocorrelation function for different child volumes
Intuitively, the sign of large child orders should have shorter memory than small ones. Traders who have large quantities $Q$ to execute is likely to trade small lots in order not to reveal information, whereas smaller $Q$ might be possible to execute in a few shots.
In order to test this hypotheses, we define five logarithmic bins for the rescaled volume $\tilde{q}=q/\phi_{D}$ where $\phi_{D}$ is the daily traded volume. We then compute for each bin the autocorrelation function $C_{\mathcal{B}(\tilde{q})}(\tau)=\operatorname{\mathbb{E}}[\varepsilon_{{B}(\tilde{q})}(t)\varepsilon_{{B}(\tilde{q})}(t+\tau)]$. We removed the largest bin, as it contains outliers, and present the four other autocorrelation functions in Fig. 3, in log-log, together with the unconditional autocorrelation function. As expected, the effective memory exponent $\gamma$ increases with $q$, going from $0.6$ to $1.3$. We take this observation as a qualitative justification of the specification proposed above, i.e. that $\mu_{q}=1+\gamma_{q}=\mu_{1}+\lambda\log q$, which as we now show, allows us to make quantitative predictions for the scaling of the generalized volume imbalances $I^{a}(T)$.
Figure 3: Evolution of the sign autocorrelation of market orders based on their corresponding volume bin $\mathcal{B}(q)$. We used data from the EUROSTOXX, between 2016 and 2019 rescaling market order volume $q$ by the daily traded volume. Our findings indicate that larger market orders tend to be less correlated than smaller ones. Note that nearly $4\%$ of market orders fall into the largest bin.
@@ 3.4.4 Scaling of the generalized volume imbalances
4 The Impact-Diffusivity puzzle and a generalized propagator
We now discuss the problem of whether the volatility of price changes can be explained chiefly in terms of the impact of intertwined metaorders of different sizes and signs that get executed in the market, i.e. without a “Fundamental” component that makes prices move without any trades. We will first recall how price diffusivity and long-memory of order flow are reconciled within the standard “propagator” framework, and then show how the issue becomes much more perplexing when the impact of metaorders obeys the square-root law. Three possible resolutions are proposed, together with their strengths and weaknesses.
4.1 Price diffusivity within the propagator model
We first assume the impact of single orders is given by a deterministic propagator model, i.e. the average price change due to an order of volume $q$ executed at time $t=0$ is $\theta(q)$ which then decays as [bouchaud2003fluctuations, bouchaud2018trades]
@@ 4.1 Price diffusivity within the propagator model
For simplicity, we assume for now that all market orders have the same volume $q$, and model the price variation $\Delta_{T}$ over time $T$ as the superposition of the average impact of metaorders, neglecting fluctuations that will be considered in section 4.5 below.
@@
Hence, we see that metaorders contribute to volatility provided $13-2\beta-\mu=$ 1. This equality coincides, as expected, with the critical condition $2\beta=1-\gamma$ derived within the propagator model (recall that $\gamma=\mu-1$). When $\beta<(1-\gamma)/2$, the price is super-diffusive (i.e. $\gg T$), whereas when $\beta>(1-\gamma)/2$, the contribution of the average impact of metaorders to price variance is negligible, i.e. $o(T)$.
In order to recover the square-root impact law, one should naively set $\beta=1/2$, such that $3-2\beta-\mu=1-\gamma$. But the contribution of metaorders to volatility $\Sigma_{T}^{2}$ would be then subdominant at long times, unless $\gamma\to 0$ (i.e. an hyper-slow decay of the sign autocorrelation function). Note that the choice $\beta=1/2^{-},\gamma=0^{+}$ corresponds to the model advocated by Jusselin & Rosenbaum [jusselin2018noarbitrageimpliespowerlawmarket], but is difficult to reconcile with the empirically determined value $\gamma\approx 0.5$ [sato2023inferring]. This value of $\gamma$, in turn, means that a square-root impact appears to be unable to generate price diffusion, since in this case $13-2\beta-\mu\approx 0.5<$ 1.
One could then argue that volatility does not primarily come from the average impact of metaorders, but rather from its fluctuations, a possibility that we explore in section 4.4 below. But in any case, the propagator model with $\beta=1/2$ fails to account for two important stylized facts:
  • Injecting $\beta=1/2$ into Eq. (42), one finds, as already mentioned above, $\mathcal{I}(s)\propto\sqrt{\widetilde{\varphi}\tau_{0}}\,\sqrt{Q}$ when $\theta(q)=\sqrt{q}$ (as indeed suggested by the data of [maitrier2025double]). Hence, one recovers the square-root law $\mathcal{I}(Q)=Y\sqrt{Q}$ but with an extra square-root dependence of the prefactor $Y$ on the participation rate $\widetilde{\varphi}$, when empirical data show that $Y$ is all but independent of $\widetilde{\varphi}$.
  • The decay of impact after the end of a metaorder, when fitted with Eq. (44), suggests a value $\beta\approx 0.2<1/2$ [bacry2015market, bucci2018slow, maitrier2025double], i.e. a much faster short time decay and a much slower long time decay than predicted by $\beta=1/2$.
@@ 4.2 A generalized propagator model
Let us assume that once a metaorder has started, the market slowly adapts and, in the spirit of the LLOB model [donier2015fullyconsistentminimalmodel, bouchaud2018trades], progressively provides more liquidity to absorb the incoming flow. We represent this effect as a two-time propagator, describing the impact of a child order occurring at time $t^{\prime}$ after the start of the metaorder on the price at time $t>t^{\prime}$:
@@
The impact of a metaorder of duration $s\gg\tau_{0}$ and $\widetilde{\varphi}s\gg n_{0}$ is now given by  [^6] Note that when $s\gg\tau_{0}$ but $\widetilde{\varphi}s\lesssim n_{0}$, impact behaves as in the standard propagator model as $t^{1-\beta}$. If $\beta\approx 0.2$, such a behaviour is much less concave than a square-root, in agreement with the results of [bucci2019crossover, maitrier2025double].
@@
which is again sub-diffusive whenever $\beta>0$ and $\gamma=\mu-1>0$. In other words, impact-driven price diffusion is only possible when impact is permanent, i.e. $\beta=0$. This is in fact the assumption made by Sato & Kanasawa in their latest paper [sato2025exactlysolvablemodelsquareroot]. However, all empirical data known to us suggest that impact decays, at least over short to medium time scales [moro2009market, brokmann2015slow, bucci2018slow, maitrier2025generatingrealisticmetaorderspublic], which according to our calculation should lead to substantial price mean reversion on such time scales.
We now turn to three possible resolutions of the diffusion “paradox”: one based on the autocorrelation of the sign of metaorders, a second one based on the permanent impact of large child orders, and the last one based on permanent impact fluctuations.
Figure 5: Comparison of the three main theories of impact decay for a metaorder of size $s=10$. During the execution phase (blue), impact follows a square-root impact growth, i.e. tproportional-$\propto\sqrt{t}$, where $t$ denotes the index of the child order [moro2009market, maitrier2025double]. In the post-execution phase (red), impact may either remain permanent (solid line) or decay (dashed). Two distinct decay mechanisms are illustrated: one based on the generalized two-time propagator as $t^{-\beta}$ with $\beta=0.2$, and the other on the LLOB theory, corresponding to $\beta=\frac{1}{2}$.
4.3 The role of metaorder autocorrelations
@@
For $\beta=0.2$, this gives $\gamma_{\times}=0.6$, which is close to $\gamma$ itself and close to its empirical value [donier2015million] and G. Maitrier (unpublished). It is also compatible with the bound obtained in Eq. (38), which ensures that the autocorrelation of signs is dominated by the size distribution of metaorders, as postulated by [lillo2005theory] and firmly established in [sato2023inferring].
@@
This result is interesting since the peak impact is then precisely given by the standard square-root law, up to a weak participation rate dependence:
@@
as indeed observed empirically (up to subleading corrections that can also be rationalized within our framework), see Fig. 6.
@@
Note that we do not observe mutifractality (i.e. $\Sigma_{T}^{(2n)}\propto T^{\zeta_{n}}$ with $\zeta_{n}\neq n$) because we work in trade time and not in real time. As it is well known, multifractal effects come from intermittent fluctuations of the activity rate $\nu$, see e.g. [muzy2000modelling, jusselin2018noarbitrageimpliespowerlawmarket]. An interesting extension of our model, which we leave for future work, would be to assume that $\nu$ itself has fractal properties.
Figure 6: Scaling of the moments of price changes $|\Delta_{T}|^{2n}$ as a function of trade time $T$. We normalized the moment values such that all curves begin at $1$ for $T=1$. As shown, a sublinear behavior is found at short times for $n>1$, likely attributable to price jumps that may dominate for short time scales when $n$ increases. Therefore, we fitted the data as $\Sigma^{(2n)}=a_{0}+a_{1}T^{\zeta_{n}}$ and present the values of $\zeta_{n}$ in the legend. We find in all cases $\zeta_{n}\approx n$, up to subleading corrections that we attribute to volume fluctuations.
4.4 The role of volume fluctuations
Another possibility is to take seriously the fact that the size $q$ of child orders is fluctuating and correlated with the duration $s$ of metaorders. As we have shown in section 3.2, a dependence of the exponent $\mu_{q}$ on $q$ explains how different moments of the volume imbalance depend on $T$, see Eqs. (32), (34). It is also in agreement with direct empirical results, see Fig. 3.
Now, similarly to $\mu_{q}$, we might assume that the impact decay exponent $\beta$ becomes $q$-dependent: $\beta_{q}=\beta_{1}-\lambda^{\prime}\log q$ [^7] A sufficient condition for price diffusivity in the standard propagator model is $\beta_{1}=1-\mu_{1}/2$ and $\lambda^{\prime}=\lambda/2$, but we will not impose such constraints below and leave $\lambda$ and $\lambda^{\prime}$ free. One then observes that price impact becomes permanent for large child orders, with $q>q_{0}$ such that $\beta_{q}\leq 0$. When metaorder signs are independent, the non-vanishing contribution to volatility then reads:
$\mathbb{E}[\Delta_{T}^{2}]_{q>q_{0}}=\nu\int_{q_{0}}^{\infty}{\rm d}q\,\Xi(q)\mathcal{I}_{1}^{2}(q,\varphi)\int_{0}^{T}{\rm d}u\,\int_{0}^{u}{\rm d}s\,\Psi_{q}(s)s\approx_{T\to\infty}\nu T\int_{q_{0}}^{\infty}{\rm d}q\,\Xi(q)\mathcal{I}_{1}^{2}(q,\varphi)\bar{s}_{q}.$ (62)
$\sigma^{2}=\frac{\mathbb{E}[\Delta_{T}^{2}]_{q>q_{0}}}{T}\propto\theta_{0}^{2}\phi_{0},$ (63)
$\theta_{0}\propto\frac{\sigma}{\sqrt{\phi_{0}}},$ (64)
$\mathcal{I}(Q|q,\widetilde{\varphi})=Y(q,\widetilde{\varphi})\sigma\sqrt{\frac{Q}{\phi}},\qquad Y(q,\widetilde{\varphi})\propto(\widetilde{\varphi}\tau_{0})^{\beta_{q}}\,\sqrt{\frac{\phi}{\phi_{0}}}.$ (65)
We see that this result suggests a weak dependence of the prefactor of the square-root law in $q$ and $\widetilde{\varphi}$, which disappears for large enough $q>q_{0}$, since $\beta_{q}\to 0$ in that case. However, in that case $Y$ would be substantially larger than empirically observed, since from Fig. 3 we estimate $\phi_{0}\approx 0.1\phi$. Besides, since child orders of size $<q_{0}$ (which are the most numerous) only impact prices temporarily, this scenario would again lead to strong visible mean-reversion in prices not observed in data – see e.g. Fig. 6 for $n=1$.
It is, however, important to discuss how volume fluctuations might affect the result Eqs. (55), (56) above, induced by the correlation between different metaorders. It is plain to extend the calculations of section 3.3 to get
$\mathbb{E}[\Delta_{T}^{2}]_{o.d.}\propto\Gamma e^{m+\frac{\sigma_{\ell}^{2}}{4}}\,T^{2-\gamma_{\times}-2\beta_{m}+\lambda^{\prime}\sigma_{\ell}^{2}},$ (66)
whereas a similar calculation for the diagonal contribution gives
$\mathbb{E}[\Delta_{T}^{2}]_{d.}\propto e^{m+\frac{\sigma_{\ell}^{2}}{2}}\,T^{1-2\beta_{m}+2\lambda^{\prime}\sigma_{\ell}^{2}}.$ (67)
4.5 The role of impact fluctuations
  • (i) news events do obviously exist (see e.g. [aubrun2025identifying] for a recent discussion) and should indeed contribute to volatility. In fact, Efficient Market Theory predicts that the only contribution to volatility comes from news!
Such a random component was in fact indirectly observed by Bucci et al. [bucci_impact_vol2019], where it was found that the effect of a single metaorder on price changes reads
$\Delta_{T}=p_{T}-p_{0}\approx\varepsilon\mathcal{I}(Q)\left[1+z\eta\right],$ (68)
where $\eta$ is a zero mean, unit variance, independent random variable, and $z$ a coefficient measuring the relative fluctuations of impact, found to be around $3$ in [bucci_impact_vol2019]. [^8] Note that there is an error in that paper, where $z$, called $a$ there, was reported to be around $0.1$.
$\Delta_{T,1}=z_{\infty}\theta_{0}\sqrt{q\varphi}\int_{0}^{T}{\rm d}N_{t}\,\varepsilon(t)\left[\mathbb{I}(t+s>T)\sqrt{T-t}+\mathbb{I}(t+s<T)\sqrt{s}\right]\eta_{t}+\sigma_{F}\xi\sqrt{T},$ (69)
$\Sigma_{T}^{2}=\mathbb{E}[\Delta_{T}^{2}]_{\eta,\xi}=z_{\infty}^{2}\theta_{0}^{2}q\varphi\nu\int_{0}^{T}{\rm d}u\int_{0}^{\infty}{\rm d}s\Psi(s)\left[\mathbb{I}(s>u)u+\mathbb{I}(s<u){s}\right]+\sigma_{F}^{2}T,$ (70)
whence a long-term volatility given by
$\sigma^{2}=\sigma_{F}^{2}+z_{\infty}^{2}\theta_{0}^{2}q\varphi\nu\bar{s}\equiv\sigma_{F}^{2}+z_{\infty}^{2}\theta_{0}^{2}\phi,$ (71)
$\sigma^{2}=\frac{\sigma_{F}^{2}}{1-Y^{2}z_{\infty}^{2}},$ (72)
i.e. excess volatility induced by trading, independently of its information content. This is in line with the empirical results of [wyart2008relation] (section 4), [taranto2018linear] and [bouchaud2018trades] (Figs. 13.2 and 14.5), and is of course related to the well-known excess volatility/excess trading puzzle, see e.g. [shiller1981stock, odean1999investors, leroy2006excess] and [bouchaud2018trades], chapters 2 & 20, see also [gabaix2021search, bouchaud2010impact, van2024ponzi] for related discussions.
$\mathcal{I}(Q)\propto\sqrt{\frac{\sigma^{2}-\sigma_{F}^{2}}{z_{\infty}^{2}}}\times\sqrt{\frac{Q}{\phi}}.$ (73)
4.6 Discussion
The second scenario, which attributes long term to the non-decaying impact of metaorders executed with large child orders does not seem very credible to us, because it would lead to substantial mean-reversion effects due to the impact decay of small child orders, which is at odds with empirical decay.
Finally, long term volatility may result from the random component of impact, assumed to be permanent. This bypasses the paradox that average impact decays and, in the absence of correlations between metaorders, should lead to subdiffusion. In such a scenario, volatility is induced by trading activity alone, even when average impact is zero. Still, the fact that impact fluctuations are proportional to average impact, as postulated in Eq. (68), is important to recover the correct relation between peak impact $\mathcal{I}(Q)$ and $\sigma\sqrt{Q/\phi}$, Eq. (73), which is now independent of $\widetilde{\varphi}$.
We now turn to the study of the covariance of price changes and volume imbalances, to see whether we can constrain the theory further. In fact, we will see that the data strongly favours an interpretation based on the average impact of long-range correlated metaorders. The last scenario, where volatility arises from the fluctuating part of impact, does not pass the test, as it is unable to reproduce the detailed structure of the correlations between price changes and generalized volume imbalances.

### v6

@@ 4.3 The role of long-range correlations between metaorders
It is known that the signs of metaorders initiated by different traders are also correlated, see [18, 15]. This may either be due to herding, or more plausibly to different traders following the same signal. As mentioned above, we assume that the sign cross-correlation $\mathbb{E}[\varepsilon(t)\varepsilon(t+\tau)]$ decays as $\Gamma(\tau_{0}/\tau)^{\gamma_{\times}}$, whereas the sizes $q$ and $q^{\prime}$ are remain independent for simplicity. [^5] One can extend the following calculations to the case where conditional size distribution of a metaorder starting at $t+\tau$, knowing that one metaorder started at $t$ is $\Psi_{q}(s^{\prime}|s,\tau)=\frac{\tau^{b}}{1+\tau^{b}}\Psi_{q}(s^{\prime})+\frac{1}{1+\tau^{b}}\frac{1}{s}F(s^{\prime}/s),$ (34) where $F(.)$ is a certain function and $b$ a new exponent. The following results are unaffected provided $b>1$. When $\Gamma=0$, there is no correlation between successive metaorders.
@@
In other words, the LMF model [34] that ascribes the main contribution to sign autocorrelation to long metaorders is only valid if such metaorders are not too strongly correlated between themselves, i.e. when
@@
In view of the empirical data supporting the LMF model, we stick to this assumption henceforth. In fact, one can measure $\gamma_{\times}$ directly (G. Maitrier, unpublished, see also [18]) suggesting $\gamma_{\times}\approx\gamma$.
@@ 4.4.1 Child volume distribution
As discussed in Section 4.2, we consider a log-normal distribution for the child order sizes, as defined by Eq. (27). It turns out to be a reasonable approximation of reality for large volumes, see Fig. 1, with values of $\sigma_{\ell}$ reported in the legend, around $1$ for stocks and SPMINI and $1.2$ for EUROSTOXX. A better fit of the tail of the distribution is, arguably, power-law $\propto q^{-1-\varpi}$, with $\varpi$ found to be around 2.12.12.42.4. Such a choice makes the mathematical analysis more cumbersome, and we prefer the log-normal specification for our semi-quantitative discussion of the role of volume fluctuations. Nevertheless, a power-law tail can play a role similar to the coefficient $\lambda$ in $\mu_{q}$ when it comes to the scaling of $\Sigma^{(2n)}_{a}$, crossing over to a linear behaviour when $2na\gtrsim\varpi$, see footnote 4.
@@ 4.4.2 Distribution of sign imbalances
We show in Fig. 2 the distribution of rescaled sign imbalances $I^{0}T^{-\chi}$, for different $T$ in trade time and $\chi=0.72$. The theoretical analysis performed in the previous sections predicts that such distributions should collapse when $\chi$ is chosen to be $1/\mu$, where $\mu=1+\gamma$ is related to the autocorrelation of the sign of the trades, which gives $\mu\approx 1.4$. Although not perfect, the agreement is quite reasonable, in view of the fact that $\mu$ may actually depend on the size of the child order $q$. The master curve is clearly non-Gaussian, with tails that become fatter as $T$ increases, as expected from our theoretical prediction.
Figure 2: Distribution of sign imbalances $I^{0}$ as a function of the window size $T$ (measured in number of trades) on EUROSTOXX, between 2016 and 2019. After a proper rescaling by $T^{-\chi}$, with $\chi=0.72$, distributions nicely collapse onto a single master curve, consistent with findings in [42]. The value of $\chi$ is not far from the theoretical prediction $\chi=1/\mu$ with $\mu=3/2$. The dotted line shows a Gaussian distribution with the same variance as red curve. As expected, the volume imbalance exhibits fat tails.
@@ 4.4.3 Scaling of the generalized volume imbalances
5 The Impact-Diffusivity puzzle and a generalized propagator model
We now discuss the problem of whether the volatility of price changes can be explained chiefly in terms of the impact of intertwined metaorders of different sizes and signs that get executed in the market, i.e. without any “fundamental” component that makes prices move without any trades. Such fundamental component would merely add a constant to the mechanical (impact induced) component computed below.
We will first recall how price diffusivity and long-memory of order flow are reconciled within the standard “propagator” framework, and then show how the issue becomes much more perplexing when the impact of metaorders obeys the square-root law. We postulate that the autocorrelation between different metaorders is responsible for the emergence of price diffusion. Two other possible resolutions are considered in Appendix A, together with their strengths and weaknesses.
5.1 Price diffusivity within the standard propagator model
We first assume the impact of single orders is given by a deterministic propagator model, i.e. the average price change due to an order of volume $q$ executed at time $t=0$ is $\theta(q)$ which then decays as [7, 5]
@@ 5.1 Price diffusivity within the standard propagator model
For simplicity, we assume for now that all market orders have the same volume $q$, and model the price variation $\Delta_{T}$ over time $T$ as the superposition of the average impact of metaorders, neglecting fluctuations that will be considered in section The role of impact fluctuations below.
@@
Hence, we see that metaorders contribute to volatility provided $3-2\beta-\mu=1$. This equality coincides, as expected, with the critical condition $2\beta=1-\gamma$ derived within the propagator model (recall that $\gamma=\mu-1$). When $\beta<(1-\gamma)/2$, the price is super-diffusive (i.e. $\gg T$), whereas when $\beta>(1-\gamma)/2$, the contribution of the average impact of metaorders to price variance is negligible, i.e. $o(T)$.
In order to recover the square-root impact law, one should naively set $\beta=1/2$, such that $3-2\beta-\mu=1-\gamma$. But the contribution of metaorders to volatility $\Sigma_{T}^{2}$ would be then subdominant at long times, unless $\gamma\to 0$ (i.e. an hyper-slow decay of the sign autocorrelation function). Note that the choice $\beta=1/2^{-},\gamma=0^{+}$ corresponds to the model advocated by Jusselin & Rosenbaum [30], but is difficult to reconcile with the empirically determined value $\gamma\approx 0.5$ [45]. This value of $\gamma$, in turn, means that a square-root impact appears to be unable to generate price diffusion, since in this case $3-2\beta-\mu\approx 0.5<1$.
One could then argue that volatility does not primarily come from the average impact of metaorders, but rather from its fluctuations, a possibility that we explore in section The role of volume fluctuations below. But in any case, the propagator model with $\beta=1/2$ fails to account for two important stylized facts:
  • Injecting $\beta=1/2$ into Eq. (41), one finds, as already mentioned above, $\mathcal{I}(s)\propto\sqrt{\widetilde{\varphi}\tau_{0}}\,\sqrt{Q}$ when $\theta(q)=\sqrt{q}$ (as indeed suggested by the data of [37]). Hence, one recovers the square-root law $\mathcal{I}(Q)=Y\sqrt{Q}$ but with an extra square-root dependence of the prefactor $Y$ on the participation rate $\widetilde{\varphi}$, when empirical data show that $Y$ is all but independent of $\widetilde{\varphi}$.
  • The decay of impact after the end of a metaorder, when fitted with Eq. (43), suggests a value $\beta\approx 0.2<1/2$ [3, 12, 37], i.e. a much faster short time decay and a much slower long time decay than predicted by $\beta=1/2$.
@@ 5.2 A generalized propagator model
Let us assume that once a metaorder has started, the market slowly adapts and, in the spirit of the LLOB model [17, 5], progressively provides more liquidity to absorb the incoming flow. We represent this effect as a two-time propagator, describing the impact of a child order occurring at time $t^{\prime}$ after the start of the metaorder on the price at time $t>t^{\prime}$:
@@
The impact of a metaorder of duration $s\gg\tau_{0}$ and $\widetilde{\varphi}s\gg n_{0}$ is now given by  [^7] Note that when $s\gg\tau_{0}$ but $\widetilde{\varphi}s\lesssim n_{0}$, impact behaves as in the standard propagator model as $t^{1-\beta}$. If $\beta\approx 0.2$, such a behaviour is much less concave than a square-root, in agreement with the results of [13, 37].
@@
which is again sub-diffusive whenever $\beta>0$ and $\gamma=\mu-1>0$. In other words, impact-driven price diffusion is only possible when impact is permanent, i.e. $\beta=0$. This is in fact the assumption made by Sato & Kanasawa in their latest paper [47]. However, all empirical data known to us suggest that impact decays, at least over short to medium time scales [39, 11, 12, 36], which according to our calculation should lead to substantial price mean reversion on such time scales.
We now discuss a possible resolution of the diffusion “paradox”, based on the autocorrelation of the sign of metaorders, which will be further validated in the next section. Two other proposals are discussed in Appendix A.
Figure 4: Comparison of the three main theories of impact decay for a metaorder of size $s=10$. During the execution phase (blue), impact follows a square-root impact growth, i.e. $\propto\sqrt{t}$, where $t$ denotes the index of the child order [39, 37]. In the post-execution phase (red), impact may either remain permanent (solid line) or decay (dashed). Two distinct decay mechanisms are illustrated: one based on the generalized two-time propagator as $t^{-\beta}$ with $\beta=0.2$, and the other on the LLOB theory, corresponding to $\beta=\frac{1}{2}$.
Figure 5: Scaling of the moments of price changes $|\Delta_{T}|^{2n}$ as a function of trade time $T$. We normalized the moment values such that all curves begin at 11 for $T=1$. As shown, a sublinear behavior is found at short times for $n>1$, likely attributable to price jumps that may dominate for short time scales when $n$ increases. Therefore, we fitted the data as $\Sigma^{(2n)}=a_{0}+a_{1}T^{\zeta_{n}}$ and present the values of $\zeta_{n}$ in the legend. We find in all cases $\zeta_{n}\approx n$, up to subleading corrections that we attribute to volume fluctuations.
5.3 Metaorder autocorrelations to the rescue
@@
For $\beta=0.2$, this gives $\gamma_{\times}=0.6$, which is close to $\gamma$ itself and close to its empirical value [18] and G. Maitrier (unpublished). It is also compatible with the bound obtained in Eq. (37), which ensures that the autocorrelation of signs is dominated by the size distribution of metaorders, as postulated by [34] and firmly established in [45].
@@
This result is interesting since the peak impact is then precisely given by the standard square-root law, up to a weak participation rate dependence when $\beta=0.2$:
@@
as indeed observed empirically (up to subleading corrections that can also be rationalized within our framework), see Fig. 5. Note that we do not observe mutifractality (i.e. $\Sigma_{T}^{(2n)}\propto T^{\zeta_{n}}$ with $\zeta_{n}\neq n$) because we work in trade time and not in real time. As it is well known, multifractal effects come from intermittent fluctuations of the activity rate $\nu$, see e.g. [40, 30]. An interesting extension of our model, which we leave for future work, would be to assume that $\nu$ itself has fractal properties.
The corrections to Eqs. (54), (55) caused by volume fluctuations and a possible dependence of $\beta$ on $q$ are further discussed in Appendix A.